OpenStudy (anonymous):

how about this one anyone y^2-6y+9=81

OpenStudy (anonymous):

\[y^2-6y+9=81\]\[y^2-6y-72=0\]\[(y-12)(y+6)=0\]\[y=12\]\[y=-6\]